# 3HDM at the ILC

### v1

@@ III.3 $h_{1}H_{2}A_{2}$ Production : $FHD$ Mode
This particular channel involving the production of $h_{1}H_{2}A_{2}$ and its subsequent decay into a fully hadronic mode with a $6b$ quark final state offers another promising avenue for exploration. The benchmark points considered here are referred to as $BP-3$ and $BP-4$ the corresponding masses, along with information regarding the branching ratios and total cross sections, are provided in Table 5.
@@
$BP-3$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow h_{1}H_{2}A_{2}$ $H_{1}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
Mass (GeV) 125 255.20 543.43 228.65 366.61 347.99 452.39 $\sigma||BR$ 0.0012179 0.54458 0.98325 0.981178 0.00063986
$BP-4$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow h_{1}H_{2}A_{2}$ $h_{1}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
Mass (GeV) 125 269.54 470.15 215.39 522.60 311.28 451.15 $\sigma||BR$ 0.001619 0.54458 0.84123 0.82652 0.000613
Table 5: The upper portion of the table represents the mass points, $BR$ and total cross section corresponding to $BP-3$, whereas the lower portion corresponds to $BP-4$.
Background Process Final State Cross Section($\sigma$) (pb)
@@
Given the highly distinctive nature of the final state, which is significantly different from typical SM backgrounds, it is reasonable to focus solely on counting the number of signal and background events based on a multiplicity cut of $N(b)=6$. At this stage, with the implementation of strong preselection cuts, it is highly unlikely that any of the SM backgrounds survive, as such a final state is relatively rare in $e^{+}e^{-}$ machines. The goal is to determine the range of luminosity required to effectively probe the production mechanism. The total cross sections for $BP-3$ and $BP-4$ are of similar magnitude, indicating that the required luminosity for discovery should also lie within a comparable range. We find that, for $BP-3$, an integrated luminosity of $1000\,\mathrm{fb}^{-1}$ and for $BP-4$, $1200\,\mathrm{fb}^{-1}$ are sufficient to achieve a discovery significance of $5\sigma$, based purely on the event count after applying the $N(b)=6$ selection. The corresponding selection cut and its effect on the event yield are summarized in the cutflow chart presented in Table 7.
$BP-3$:Cuts Signal (S) Background (B) S vs B
@@
$BP-4$:Cuts Signal (S) Background (B) S vs B
Initial (no cut) 735 110509
SEL: $N(b)=6$ 32.28 6.0 5.217
Table 7: The upper portion of the table corresponds to the cutflow chart for $BP-3$ at $\mathcal{L}=1000\,fb^{-1}$, while the lower portion corresponds to $BP-4$ at $\mathcal{L}=1200\,fb^{-1}$.
@@ III.4 $H_{2}H_{2}Z$ Production : $FHD$ Mode
Although the signal contains exactly four $b$-quarks at the parton level, detector level reconstruction is impacted by $b$-tagging inefficiencies and possible contamination from gluon splitting processes. Accordingly, a requirement of $N(b)\geq 4$ is imposed to retain signal efficiency while suppressing background contributions, and is adopted along with a cut of $N(j)\geq 2$. Following the structure of the previously discussed channels, we consider two benchmark points for the analysis, denoted as $BP\text{-}5$ and $BP\text{-}6$, which are consistent with current experimental constraints. The relevant SM backgrounds for this process are listed in Table 8.555This same background table is also referenced in the $A_{2}A_{2}Z$ production case, as both processes in the FHD mode yield the same final state of $4b+2j$. Detailed information on the chosen benchmark points—including their masses, branching ratios, and total production cross sections—is presented in Table 9.
Background Process Final State Cross Section($\sigma$) (pb)
@@
Table 8: The relevant SM backgrounds for the fully hadronic decay mode of $H_{2}H_{2}Z$ (Section III.4) and $A_{2}A_{2}Z$ (Section III.5) production.
BP-5:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 207.83 527.73 315.17 397.45 333.29 417.01 $\sigma||BR$ 0.007175 0.815618 0.815618 0.548 0.002616
BP-6:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 219.73 418.47 339.33 398.26 351.98 468.76 $\sigma||BR$ 0.006474 0.868647 0.868647 0.548 0.002677
Table 9: The upper table corresponds to $BP-5$, and the lower table to $BP-6$, summarizing the benchmark mass points along with the associated $BR$ and total cross sections.
The selection cuts $N(b)\geq 4$ and $N(j)\geq 2$ are employed to preferentially isolate events consistent with the signal topology, thereby suppressing contributions from SM backgrounds. This approach, while simple in implementation, proves effective in enhancing the signal purity of the sample. Nevertheless, an important consequence of these inclusive number-based cuts is the significant overlap observed in the distributions of key kinematic observables between signal and background as can be evidenced from Fig. 3. As a result, no single observable emerges as a strong discriminator suitable for additional shape-based selection. Despite this limitation, the overall event yield remains sufficiently high. At an integrated luminosity of $300\,\mathrm{fb}^{-1}$, the signal reaches the threshold for discovery with a statistical significance of $5\sigma$. The cut-based selection criteria data is provided in Table 10.
@@
$BP-5$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 1000 30272
SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 95.92 247.0 5.180
$BP-6$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 1024 30272
SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 112.9 247.0 5.951
Table 10: The upper panel presents the cutflow summary for the BP-5 signal, while the lower panel shows the corresponding results for the BP-6 signal. Both cases assume an integrated luminosity of $300\,\mathrm{fb}^{-1}$.
@@ III.5 $A_{2}A_{2}Z$ Production : $FHD$ Mode
The $A_{2}A_{2}Z$ production channel and its subsequent decay topology are identical to the $H_{2}H_{2}Z$ case discussed in Section III.4. The corresponding SM backgrounds are listed in Table 8. The signal benchmarks, labeled as $BP\text{-}7$ and $BP\text{-}8$, correspond to different mass configurations. Detailed information on their mass values, BR’s and total production cross sections is provided in Table 11. The signal process is given by:
@@
A similar strategy is adopted for signal isolation in this channel, guided by the final state topology of $4b+2j$. In particular, number-based selection criteria of $N(b)\geq 4$ and $N(j)\geq 2$ are applied, consistent with the methodology outlined in the previous section. The cutflow summary for benchmark points $BP\text{-}7$ and $BP\text{-}8$ is provided in Table 12. Notably, even with only these inclusive number cuts, a favorable balance between signal efficiency and background suppression is achieved, highlighting the effectiveness of this minimal selection strategy.
BP-7:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 354.97 552.52 210.87 391.04 369.26 404.86 $\sigma||BR$ 0.00637189 0.8188199 0.8188199 0.548 0.002341
BP-8:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 354.62 385.08 202.86 548.99 322.35 551.64 $\sigma||BR$ 0.00661869 0.864266 0.864266 0.548 0.002709
Table 11: The upper table represents the mass points, BR and total cross section information for $BP-7$, whereas the lower table represents the same for $BP-8$.
$BP-7$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 895 30272
SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 108.43 247.0 5.752
$BP-8$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 1036 30272
SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 123.2 247.0 6.402
Table 12: The upper panel presents the cutflow summary for the BP-7 signal, while the lower panel shows the corresponding results for the BP-8 signal. Both cases are analyzed at an integrated luminosity of $300\,\mathrm{fb}^{-1}$.
@@ III.6 $H_{2}H_{2}Z$ Production : $SLD$ Mode
The $H_{2}H_{2}Z$ production channel discussed earlier in Section III.4 can also be considered with an alternative decay mode, wherein the $Z$ boson decays leptonically. The signal process is thus given by:
@@
This leads to a distinct final state of $4b+2\ell$. One of the consequences could be a drop in the total cross section, as the $Z$ decaying hadronically yields a higher branching ratio, whereas its leptonic decay significantly reduces the total cross section due to a much lower branching ratio. Nevertheless, it must be compared against distinct and appropriate SM backgrounds. To proceed with the analysis, it is necessary to identify Standard Model backgrounds that can mimic this signature. The relevant backgrounds are listed in Table 13666The background processes listed apply equally to both $H_{2}H_{2}Z$ and $A_{2}A_{2}Z$ production channels in the semi-leptonic decay mode, as they yield identical final states.. Details of the two benchmark points, $BP\text{-}9$ and $BP\text{-}10$, including their masses, branching ratios, and cross sections, are provided in Table 14.
Background Process Final State Cross Section($\sigma$) (pb)
@@
Table 13: The relevant SM backgrounds for the semi-leptonic decay mode of $H_{2}H_{2}Z$ (Section III.6) and $A_{2}A_{2}Z$ (Section III.7) channels. The demand of at least two leptons necessitates the presence of SM gauge bosons that might also decay to neutrinos, and thus many of these backgrounds have sizable missing energy.
BP-9:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
@@
BP-10:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
Mass (GeV) 125 219.73 418.47 339.33 398.26 351.98 468.76 $\sigma||BR$ 0.006474 0.868647 0.868647 0.0673 0.0003288
Table 14: The upper portion of the table represents the mass points, BR, and total cross-section information for $BP-9$, while the lower portion represents the same for $BP-10$.
The analysis begins by applying number-based selection criteria, requiring $N(b)\geq 4$ and $N(\ell)\geq 2$, to facilitate effective reconstruction of the signal. Here, in this case as well, once the required number cuts are implemented, the remaining observable distributions for the signal and backgrounds become highly overlapping, making it difficult to distinguish between them. Also given that the BR for leptonic decays is significantly lower than that of the hadronic mode discussed earlier, a higher integrated luminosity is generally required to achieve comparable signal sensitivity. The cutflow summary for benchmark points $BP\text{-}9$ and $BP\text{-}10$ is presented in Table 15.
$BP-9$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 954 22147
SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 40.85 17.04 5.368
$BP-10$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 976 22147
SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 49.23 17.04 6.05
Table 15: The upper panel presents the cutflow summary for the BP-9 signal, while the lower panel shows the corresponding results for the BP-10 signal. Both cases assume an integrated luminosity of $2000\,\mathrm{fb}^{-1}$.

### v2

@@ III.3 $h_{1}H_{2}A_{2}$ Production : $FHD$ Mode
This particular channel involving the production of $h_{1}H_{2}A_{2}$ and its subsequent decay into a fully hadronic mode with a $6b$ quark final state offers another promising avenue for exploration. The benchmark point considered here is referred to as BP2 - the corresponding masses, along with information regarding the branching ratios and total cross sections, are provided in Table 5.
@@
BP2:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow h_{1}H_{2}A_{2}$ $h_{1}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
Mass (GeV) 125 255.20 543.43 228.65 366.61 347.99 452.39 $\sigma\times BR$ 0.0012179 0.54458 0.98325 0.981178 0.00063986
Table 5: The upper portion of the table represents the mass points, BR and total cross section corresponding to BP2.
Background Process Final State Cross Section ($\sigma$) [pb]
@@
Given the highly distinctive nature of the final state, which is significantly different from typical SM backgrounds, it is reasonable to focus solely on counting the number of signal and background events based on a multiplicity cut of $N(b)=6$. At this stage, with the implementation of strong preselection cuts, it is highly unlikely that any of the SM backgrounds survive, as such a final state is relatively rare in $e^{+}e^{-}$ machines. The goal is to determine the range of luminosity required to effectively probe the production mechanism. We find that, for BP2, an integrated luminosity of $1000\,\mathrm{fb}^{-1}$ are sufficient to achieve a discovery significance of $5\sigma$, based purely on the event count after applying the $N(b)=6$ selection. The corresponding selection cut and its effect on the event yield are summarized in the cutflow chart presented in Table 7.
BP2 : Cuts Signal (S) Background (B) S vs B
@@
Table 7: The upper portion of the table corresponds to the cutflow chart for BP3 at $\mathcal{L}=1000\,fb^{-1}$, while the lower portion corresponds to BP4 at $\mathcal{L}=1200\,fb^{-1}$.
@@ III.4 $H_{2}H_{2}Z$ Production : $FHD$ Mode
Although the signal contains exactly four $b$-quarks at the parton level, detector level reconstruction is impacted by $b$-tagging inefficiencies and possible contamination from gluon splitting processes. Accordingly, a requirement of $N(b)\geq 4$ is imposed to retain signal efficiency while suppressing background contributions, and is adopted along with a cut of $N(j)\geq 2$. Following the structure of the previously discussed channels, we consider a benchmark point for the analysis, denoted as BP3, which is consistent with current experimental constraints. The relevant SM backgrounds for this process are listed in Table 8.666This same background table is also referenced in the $A_{2}A_{2}Z$ production case, as both processes in the FHD mode yield the same final state of $4b+2j$. Detailed information on the chosen benchmark points—including their masses, branching ratios, and total production cross sections—is presented in Table 9.
Background Process Final State Cross Section ($\sigma$) [pb]
@@
Table 8: The relevant SM backgrounds for the fully hadronic decay mode of $H_{2}H_{2}Z$ (subsection III.4) and $A_{2}A_{2}Z$ (subsection III.5) production.
BP3:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 207.83 527.73 315.17 397.45 333.29 417.01 $\sigma\times BR$ 0.007175 0.815618 0.815618 0.548 0.002616
Table 9: The table corresponds to BP3, summarizing the benchmark mass points along with the associated BR and total cross sections.
The selection cuts $N(b)\geq 4$ and $N(j)\geq 2$ are employed to preferentially isolate events consistent with the signal topology, thereby suppressing contributions from SM backgrounds. This approach, while simple in implementation, proves effective in enhancing the signal purity of the sample. Nevertheless, an important consequence of these inclusive number-based cuts is the significant overlap observed in the distributions of key kinematic observables between signal and background as can be evidenced from Figure 3. As a result, no single observable emerges as a strong discriminator suitable for additional shape-based selection. Despite this limitation, the overall event yield remains sufficiently high. At an integrated luminosity of $500\,\mathrm{fb}^{-1}$, the signal reaches the threshold for discovery with a statistical significance of $5\sigma$. The cut-based selection criteria data is provided in Table 10.
@@
BP3 : Cuts Signal (S) Background (B) S vs B
Initial (no cut) 1307 50454
SEL: $N(b)\geq 4$ & $N(j)\geq 2$ 125.4 411.6 5.410
Table 10: The panel presents the cutflow summary for the BP3 signal. Here we assume an integrated luminosity of $500\,\mathrm{fb}^{-1}$.
@@ III.5 $A_{2}A_{2}Z$ Production : $FHD$ Mode
The $A_{2}A_{2}Z$ production channel and its subsequent decay topology are identical to the $H_{2}H_{2}Z$ case discussed in subsection III.4. The corresponding SM backgrounds are listed in Table 8. The signal benchmarks, labeled as BP7 and BP8, correspond to different mass configurations. Detailed information on their mass values, BR’s and total production cross sections is provided in Table 11. The signal process is given by:
@@
A similar strategy is adopted for signal isolation in this channel, guided by the final state topology of $4b+2j$. In particular, number-based selection criteria of $N(b)\geq 4$ and $N(j)\geq 2$ are applied, consistent with the methodology outlined in the previous section. The cutflow summary for benchmark points BP4 is provided in Table 12. Notably, even with only these inclusive number cuts, a favorable balance between signal efficiency and background suppression is achieved, highlighting the effectiveness of this minimal selection strategy.
BP4:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 354.97 552.52 210.87 391.04 369.26 404.86 $\sigma\times BR$ 0.00637189 0.8188199 0.8188199 0.548 0.002341
Table 11: The table represents the mass points, BR and total cross section information for BP4.
BP4 : Cuts Signal (S) Background (B) S vs B
Initial (no cut) 936 40363
SEL: $N(b)\geq 4$ & $N(j)\geq 2$ 113.42 329.3 5.391
Table 12: The upper panel presents the cutflow summary for the BP4 signal. This case is analyzed at an integrated luminosity of $400\,\mathrm{fb}^{-1}$.
@@ III.6 $H_{2}H_{2}Z$ Production : $SLD$ Mode
The $H_{2}H_{2}Z$ production channel discussed earlier in subsection III.4 can also be considered with an alternative decay mode, wherein the $Z$ boson decays leptonically. The signal process is thus given by:
@@
This leads to a distinct final state of $4b+2\ell$. One of the consequences could be a drop in the total cross section, as the $Z$ decaying hadronically yields a higher branching ratio, whereas its leptonic decay significantly reduces the total cross section due to a much lower branching ratio. Nevertheless, it must be compared against distinct and appropriate SM backgrounds. To proceed with the analysis, it is necessary to identify Standard Model backgrounds that can mimic this signature. The relevant backgrounds are listed in Table 13777The background processes listed apply equally to both $H_{2}H_{2}Z$ and $A_{2}A_{2}Z$ production channels in the semi-leptonic decay mode, as they yield identical final states.. Details of the benchmark point, BP5, including their masses, branching ratios, and cross sections, are provided in Table 14.
Background Process Final State Cross Section $\sigma$ (pb)
@@
Table 13: The relevant SM backgrounds for the semi-leptonic decay mode of $H_{2}H_{2}Z$ (subsection III.6) and $A_{2}A_{2}Z$ (subsection III.7) channels. The demand of at least two leptons necessitates the presence of SM gauge bosons that might also decay to neutrinos, and thus many of these backgrounds have sizable missing energy.
BP5:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
@@
Table 14: The table represents the mass points, BR, and total cross-section information for BP5.
The analysis begins by applying number-based selection criteria, requiring $N(b)\geq 4$ and $N(\ell)\geq 2$, to facilitate effective reconstruction of the signal. Here, in this case as well, once the required number cuts are implemented, the remaining observable distributions for the signal and backgrounds become highly overlapping, making it difficult to distinguish between them. Also given that the BR for leptonic decays is significantly lower than that of the hadronic mode discussed earlier, a higher integrated luminosity is generally required to achieve comparable signal sensitivity. The cutflow summary for benchmark points $BP5$ is presented in Table 15.
BP5 : Cuts Signal (S) Background (B) S vs B
Initial (no cut) 963 33221
SEL: $N(b)\geq 4$ &\& $N(\ell)\geq 2$ 41.26 25.6 5.047
Table 15: The panel presents the cutflow summary for the BP5 signal. Here we assume an integrated luminosity of $3000\,\mathrm{fb}^{-1}$.